Find a third degree McLaurin expansion of f (x) =sec2x
1
Expert's answer
2020-05-21T12:56:20-0400

Consider the function "f(x)=sec(2x)"

The McLaurin series of a function "f(x)" is given by the expansion formula as follows:


"f(x)\approx f(0)+\frac{f'(0)}{1!}x+\frac{f''(0)}{2!}x^2+\frac{f'''(0)}{3!}x^3+...+..."


Now, find the first derivative of the function as,


"f'(x)=2sec(2x)tan(2x)"

Find the second derivative of the function as,


"f''(x)=2[2sec(2x)tan(2x)tan(2x)+sec(2x)(2sec^2(2x))]"

"f''(x)=4sec(2x)tan^2(2x)+4sec^3(2x)"

Find the third derivative of the function as,


"f'''(x)=4[sec(2x)\frac{d}{dx}tan^2(2x)+tan^2(2x)\frac{d}{dx}sec(2x)]+4[\frac{d}{dx}sec^3(2x)]"


"f'''(x)=4[2tan(2x)sec^3(2x)+2tan^3(2x)sec(2x)]+4[6sec^3(2x)tan(2x)]"

Find the fourth derivative of the function as,

"f^{iv}(x)=2[(-4\left(-2\sec \left(2x\right)+4\sec ^3\left(2x\right)\right)+24\left(6\sec ^3\left(2x\right)\tan ^2\left(2x\right)+2\sec ^5\left(2x\right)\right)]"


Next, find the function value and derivative value at "x=0" as,


"f(0)=1"


"f'(0)=0"


"f''(0)=4"


"f'''(0)=0"


"f^{iv}(0)=80"


So, the third order McLaurin expansion is given by,


"f(x)\approx 1+\frac{0}{1}x+\frac{4}{2}x^2+\frac{0}{6}x^3+\frac{80}{24}x^4+..."


"f(x)\approx 1+2x^2+0\cdot x^3+..."


Therefore, the third degree McLaurin expansion of the function "f(x)=sec(2x)" is "f(x)\approx 1+2x^2+0\cdot x^3"